# Memory-Dominated Quantum Criticality as a Universal Route to High-Temperature Superconductivity

### v1

@@ Abstract
$\tilde{X}_{\lambda}(t)$ $=g(\lambda)\int dt^{\prime}\;G_{\lambda}^{A}(t-t^{\prime})\,\tilde{x}(t^{\prime}).$ (S25)
Substituting (S25) into (S21) eliminates $\tilde{X}_{\lambda}$ entirely and yields an effective contribution to the (x,x~)($\tilde{x}$) action.
@@
1. The $\tilde{X}_{\lambda}x$ term (deterministic memory kernel).
@@
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t).$ (S26)
@@
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)=\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,\tilde{x}(t^{\prime})\,x(t).$ (S27)
@@
$\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t^{\prime}-t)\,\tilde{x}(t)\,x(t^{\prime}).$ (S28)
@@
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,x(t)$ $=\int dt\,dt^{\prime}\;\tilde{x}(t)\,\underbrace{g(\lambda)^{2}\,G_{\lambda}^{R}(t-t^{\prime})}_{\equiv K_{\lambda}(t-t^{\prime})}\,x(t^{\prime}).$ (S30)
@@
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}$ (S33)
$=\int dt\,dt^{\prime}\,dt^{\prime\prime}\;D_{\lambda}\,g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,G_{\lambda}^{A}(t-t^{\prime\prime})\,\tilde{x}(t^{\prime})\,\tilde{x}(t^{\prime\prime}).$ (S34)
@@
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}=\int dt^{\prime}\,dt^{\prime\prime}\;\tilde{x}(t^{\prime})\,\mathcal{N}_{\lambda}(t^{\prime},t^{\prime\prime})\,\tilde{x}(t^{\prime\prime}).$ (S36)
@@
$S_{\mathrm{eff}}[x,\tilde{x}]=\int dt\;\tilde{x}(t)\big(\dot{x}(t)+rx(t)\big)-\int dt\;D_{x}\,\tilde{x}(t)^{2}$
$-\int dt\,dt^{\prime}\;\tilde{x}(t)\,K(t-t^{\prime})\,x(t^{\prime})-\int dt\,dt^{\prime}\;\tilde{x}(t)\,\mathcal{N}(t,t^{\prime})\,\tilde{x}(t^{\prime}),$ (S37)
@@
$\int dt\,dt^{\prime}\;\tilde{x}(t)\,K(t-t^{\prime})\,x(t^{\prime})$ $=\int\frac{d\omega}{2\pi}\;\tilde{x}(-\omega)\,K(\omega)\,x(\omega),$ (S40)
@@
In the MSRJD response sector, the retarded inverse propagator for $x$ is read off from the $\tilde{x}(-\omega)x(\omega)$ coefficient in $S_{\mathrm{eff}}$:
@@
In our framework the natural object is the time-scale density of states $\rho(\lambda)$ of relaxation rates (e.g. from Jacobian eigenvalues). In the simplest case $\overline{g^{2}}(\lambda)$ varies weakly over the slow sector so that $\mathcal{J}(\lambda)\propto\rho(\lambda)$ in the infrared, and therefore
@@
This makes explicit, at the level of a controlled Gaussian MSRJD integration, why the relaxation-rate spectrum provides the natural basis for the effective dynamics and response of the collective variable.
@@
Appendix B Appendix B: Exact emergence of a memory kernel from integrating out an OU reservoir
In this Appendix we provide a self-contained derivation showing that integrating out a continuum of relaxational Ornstein–Uhlenbeck (OU) modes generates a nonlocal-in-time (memory) kernel for the remaining collective coordinate. The result makes explicit that long-memory response is the generic outcome of eliminating hidden internal degrees of freedom, and that a flat time-scale density of states yields the universal tail $K(t)\sim 1/t$.
(I) Coupled Markovian dynamics: collective field plus relaxational reservoir: For clarity we first suppress spatial dependence and focus on temporal structure; the extension to $\mathbf{r}$ (including $cq^{2}$ terms) is immediate by Fourier transforming in space. We consider a slow collective variable $\phi(t)$ linearly coupled to a continuum of relaxational OU modes $X_{\lambda}(t)$ labeled by decay rate $\lambda\geq 0$:
@@ Appendix B Appendix B: Exact emergence of a memory kernel from integrating out an OU reservoir
Here $r>0$ is the bare restoring (mass) parameter, $\Lambda$ is an ultraviolet cutoff on relaxation rates, and $\eta$ and $\zeta_{\lambda}$ are (possibly independent) noise sources. Equations (S54)–(S55) define a Markovian (first-order) dynamical system in the enlarged state space $\{\phi,X_{\lambda}\}$.
@@
$X_{\lambda}(t)$ (S59)
$\quad=e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})+\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\Big[g(\lambda)\phi(s)+\zeta_{\lambda}(s)\Big].$
@@
$=-r\,\phi(t)+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\,e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})$
$\qquad+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)$
$\qquad+\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\zeta_{\lambda}(s)+\eta(t).$ (S60)
We now isolate the term proportional to the history $\phi(s)$:
$\int_{0}^{\Lambda}d\lambda\,g(\lambda)\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}g(\lambda)\phi(s)$
$\quad=\int_{t_{0}}^{t}ds\,\Bigg[\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}e^{-\lambda(t-s)}\Bigg]\phi(s).$ (S61)
This motivates the definition of the (causal) memory kernel
$K(t-s)\equiv\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\,e^{-\lambda(t-s)},\qquad t\geq s,$ (S62)
so that (S61) becomes simply
@@
With this definition, Eq. (S60) can be written as
@@
For $t-t_{0}\gg\Lambda^{-1}$, the transient $J_{\rm ic}(t)$ decays. More precisely, decay holds provided the effective spectral weight $g(\lambda)X_{\lambda}(t_{0})$ is integrable near $\lambda=0$ so that the $\lambda\to 0$ sector does not produce a non-decaying contribution.
(IV) TDOS form and Laplace-transform structure: Defining the time-scale density of states as
$\rho(\lambda)\equiv g(\lambda)^{2},$ (S67)
(Here $\rho(\lambda)$ should be understood as an effective spectral weight for the coupling of relaxation rates to $\phi$; in a microscopic discrete-mode representation one naturally has $\rho(\lambda)\overline{g^{2}}(\lambda)$, and we absorb any slowly varying $\overline{g^{2}}(\lambda)$ into $\rho(\lambda)$ for notational simplicity.)
the memory kernel (S62) becomes
@@
Thus $K(t)$ is precisely the Laplace transform of the TDOS (with cutoff $\Lambda$). In this sense the TDOS is the fundamental spectral object, and the time-domain memory kernel is its direct transform.
@@
$K(t)$ $=\int_{0}^{\Lambda}d\lambda\,\rho_{0}\,e^{-\lambda t}=\rho_{0}\int_{0}^{\Lambda}d\lambda\,e^{-\lambda t}.$ (S70)
@@
$K(t)$ $=\rho_{0}\int_{0}^{\Lambda t}\frac{du}{t}\,e^{-u}=\frac{\rho_{0}}{t}\int_{0}^{\Lambda t}du\,e^{-u}=\frac{\rho_{0}}{t}\Big(1-e^{-\Lambda t}\Big).$ (S71)
@@
(We use the convention $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma^{R}(\omega)$; other sign conventions for the self-energy appear in the literature.)
@@
(VII) Conceptual structure: from hidden modes to infrared memory: The derivation above establishes a general structural result. A Markovian dynamical system defined in an extended state space $\{\phi,X_{\lambda}\}$ becomes non-Markovian when the internal degrees of freedom $X_{\lambda}$ are eliminated. The resulting equation for $\phi$ necessarily acquires a causal memory kernel given exactly by the Laplace transform of the time-scale density of states (TDOS),
@@
$K(t)\simeq\frac{\rho_{0}}{t},\qquad t\gg\Lambda^{-1}.$ (S83)
@@
Appendix C Appendix C: Renormalization-group flow of the Cooper channel in conventional and memory-dominated regimes
@@
Appendix D Appendix D: Finite-temperature regulation of the infrared divergence
@@
Appendix E Appendix E: Spectral function generated by memory–dominated self–energy

### v2

@@ Appendix A Appendix A: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
$\tilde{X}_{\lambda}(t)$ $=g(\lambda)\int dt^{\prime}\;G_{\lambda}^{A}(t-t^{\prime})\,\tilde{\phi}(t^{\prime}).$ (S41)
[…]
@@
1. The $\tilde{X}_{\lambda}\phi$ term (deterministic memory kernel).
@@
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t).$ (S42)
@@
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)=\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,\tilde{\phi}(t^{\prime})\,\phi(t).$ (S43)
@@
$\int dt\,dt^{\prime}\;g(\lambda)^{2}\,G_{\lambda}^{A}(t^{\prime}-t)\,\tilde{\phi}(t)\,\phi(t^{\prime}).$ (S44)
@@
$\int dt\;g(\lambda)\,\tilde{X}_{\lambda}(t)\,\phi(t)$ $=\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,\underbrace{g(\lambda)^{2}\,G_{\lambda}^{R}(t-t^{\prime})}_{\equiv K_{\lambda}(t-t^{\prime})}\,\phi(t^{\prime}).$ (S46)
@@
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}$
$=\int dt\,dt^{\prime}\,dt^{\prime\prime}\;D_{\lambda}\,g(\lambda)^{2}\,G_{\lambda}^{A}(t-t^{\prime})\,G_{\lambda}^{A}(t-t^{\prime\prime})\,\tilde{\phi}(t^{\prime})\,\tilde{\phi}(t^{\prime\prime}).$ (S49)
@@
$\int dt\;D_{\lambda}\,\tilde{X}_{\lambda}(t)^{2}=\int dt^{\prime}\,dt^{\prime\prime}\;\tilde{\phi}(t^{\prime})\,\mathcal{N}_{\lambda}(t^{\prime},t^{\prime\prime})\,\tilde{\phi}(t^{\prime\prime}).$ (S51)
@@
$S_{\mathrm{eff}}[\phi,\tilde{\phi}]=\int dt\;\tilde{\phi}(t)\big(\dot{\phi}(t)+r\phi(t)\big)-\int dt\;D_{\phi}\,\tilde{\phi}(t)^{2}$
$-\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,K(t-t^{\prime})\,\phi(t^{\prime})-\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,\mathcal{N}(t,t^{\prime})\,\tilde{\phi}(t^{\prime}),$ (S52)
@@
$\int dt\,dt^{\prime}\;\tilde{\phi}(t)\,K(t-t^{\prime})\,\phi(t^{\prime})$ $=\int\frac{d\omega}{2\pi}\;\tilde{\phi}(-\omega)\,K(\omega)\,\phi(\omega),$ (S55)
@@
In the MSRJD response sector, the retarded inverse propagator for $\phi$ is read off from the $\tilde{\phi}(-\omega)\phi(\omega)$ coefficient in $S_{\mathrm{eff}}$:
@@
In our framework the natural object is the time-scale density of states $\rho(\lambda)$ of relaxation rates. In the simplest case $\overline{g^{2}}(\lambda)$ varies weakly over the slow sector so that $\mathcal{J}(\lambda)\propto\rho(\lambda)$ in the infrared, and therefore
@@
This makes explicit, at the level of a controlled Gaussian MSRJD integration, why the relaxation-rate spectrum provides the natural basis for the effective dynamics and response of the collective field.
@@
(X) Relation to the Keldysh functional integral: We briefly clarify the formal relation between the MSRJD formulation employed in the present work and the real–time Keldysh functional integral for driven–dissipative quantum systems.
The MSRJD action for a stochastic dynamical variable $\phi(t)$ with memory kernel and Gaussian noise may be written as
$S_{\mathrm{MSRJD}}$ $=\int dt\,dt^{\prime}\,\tilde{\phi}(t)\left[\delta(t-t^{\prime})(\partial_{t}+r)-K_{R}(t-t^{\prime})\right]$ (S69)
$\quad\phi(t^{\prime})-\int dt\,dt^{\prime}\,\tilde{\phi}(t)\,\mathcal{N}(t,t^{\prime})\,\tilde{\phi}(t^{\prime}),$
where $\tilde{\phi}$ denotes the response field, $K_{R}$ is the retarded dynamical kernel (memory kernel), and $\mathcal{N}$ is the noise correlator introduced above.
The real–time Keldysh functional integral for an open quantum system is naturally formulated in terms of classical and quantum fields, $\phi_{\mathrm{cl}}$ and $\phi_{q}$, defined by the Keldysh rotation
$\phi_{\mathrm{cl}}=\frac{\phi_{+}+\phi_{-}}{2},\qquad\phi_{q}=\phi_{+}-\phi_{-},$ (S70)
where $\phi_{\pm}$ denote fields on the forward and backward branches of the time contour.
To quadratic order in the quantum field, the Keldysh action assumes the general structure
$S_{\mathrm{K}}=\int_{\omega}\begin{pmatrix}\phi_{\mathrm{cl}}&\phi_{q}\end{pmatrix}_{-\omega}\begin{pmatrix}0&K_{A}(\omega)\\ K_{R}(\omega)&iK_{K}(\omega)\end{pmatrix}\begin{pmatrix}\phi_{\mathrm{cl}}\\ \phi_{q}\end{pmatrix}_{\omega},$ (S71)
where $K_{R,A}$ are retarded/advanced kernels and $K_{K}$ is the Keldysh (noise) component. The retarded kernel encodes dissipative response, while the Keldysh component determines fluctuation amplitudes.
In the regime of strong decoherence or large occupation, where quantum interference effects are suppressed, the quantum field $\phi_{q}$ acts as a response field. Identifying
$\phi_{q}\;\leftrightarrow\;i\tilde{\phi}$, (S72)
and integrating out $\phi_{q}$ via a Hubbard–Stratonovich transformation, one recovers a stochastic Langevin equation with noise correlator set by $K_{K}$. The corresponding MSRJD action then takes the form (S69), with the correspondence
$K_{R}(\omega)\;\leftrightarrow\;K_{R}(\omega),\qquad\mathcal{N}(\omega)\;\leftrightarrow\;\tfrac{1}{2}K_{K}(\omega).$ (S73)
Thus the MSRJD formulation used in the present work can be viewed as the classical (or strongly dissipative) limit of the Keldysh functional integral, in which the relevant dynamical information is encoded in the retarded kernel and its associated spectral properties.
Importantly, our theory does not rely on coherent quantum interference or operator-valued field structure. Instead, the central object is the spectrum of the dynamical generator (equivalently, the retarded kernel), whose low-frequency density of states controls the infrared behavior and pairing amplification mechanism discussed in the main text.
Appendix C Appendix C: Exact emergence of a memory kernel from integrating out an OU reservoir
In this Appendix we provide a self-contained derivation showing that integrating out a continuum of relaxational Ornstein–Uhlenbeck (OU) modes generates a nonlocal-in-time (memory) kernel for the remaining collective coordinate. The result makes explicit that long-memory response is the generic outcome of eliminating hidden internal degrees of freedom, and that a flat time-scale density of states yields the universal tail $K(t)\sim 1/t$. This time-domain elimination is equivalent in content to the Gaussian MSRJD integration, but is presented here directly at the level of the coupled equations of motion.
(I) Coupled Markovian dynamics (collective field plus relaxational reservoir): For clarity we first suppress spatial dependence and focus on temporal structure; the extension to $\mathbf{r}$ (including $cq^{2}$ terms) is immediate by Fourier transforming in space. We consider a slow collective variable $\phi(t)$ linearly coupled to a continuum of relaxational OU modes $X_{\lambda}(t)$ labeled by decay rate $\lambda\geq 0$:
@@ Appendix C Appendix C: Exact emergence of a memory kernel from integrating out an OU reservoir
Here $r>0$ is the bare restoring (mass) parameter, $\Lambda$ is an ultraviolet cutoff on relaxation rates, and $\eta$ and $\zeta_{\lambda}$ are (possibly independent) noise sources. For $\lambda>0$ each reservoir coordinate is exponentially stable; the sector $\lambda\to 0^{+}$ should be understood as the infrared limit of a continuous spectrum of increasingly slow relaxational modes rather than an exactly conserved degree of freedom. Equations (S74)–(S75) define a Markovian (first-order) dynamical system in the enlarged state space $\{\phi,X_{\lambda}\}$.
@@
$X_{\lambda}(t)$ $=e^{-\lambda(t-t_{0})}X_{\lambda}(t_{0})$ (S79)
$\quad+\int_{t_{0}}^{t}ds\,e^{-\lambda(t-s)}\Big[g(\lambda)\phi(s)+\zeta_{\lambda}(s)\Big].$
@@
This motivates the definition of the causal memory kernel
$K(t-s)\equiv\Theta(t-s)\int_{0}^{\Lambda}d\lambda\,g(\lambda)^{2}\,e^{-\lambda(t-s)},$ (S81)
so that (LABEL:eq:kernel_identification_step_app) becomes
@@
With this definition, Eq. (LABEL:eq:phi_substituted_app) can be written as
@@
For $t-t_{0}\gg\Lambda^{-1}$, the transient $J_{\rm ic}(t)$ decays. More precisely, decay holds provided the effective spectral weight $g(\lambda)X_{\lambda}(t_{0})$ is integrable near $\lambda=0$ so that the $\lambda\to 0^{+}$ sector does not produce a non-decaying contribution.
(IV) TDOS form and Laplace-transform structure: The memory kernel identified above depends only on the combination $g(\lambda)^{2}$ that weights the contribution of each relaxational mode to the collective dynamics. It is therefore natural to introduce the TDOS as the effective coupling-weighted spectral density of relaxation rates:
$\rho(\lambda)\equiv g(\lambda)^{2}.$ (S86)
More generally, in a microscopic description with discrete relaxational modes of density $\nu(\lambda)$ and coupling amplitudes $g_{\lambda}$, the kernel involves the combination $\nu(\lambda)|g_{\lambda}|^{2}$. For notational simplicity, we absorb any slowly varying factors into the definition of $\rho(\lambda)$. Thus $\rho(\lambda)$ should be understood as the effective spectral weight governing how strongly relaxation modes at scale $\lambda$ influence the collective coordinate $\phi$.
With this definition, the memory kernel becomes
@@
Equation (S87) shows that the memory kernel is precisely the Laplace transform of the TDOS (with ultraviolet cutoff $\Lambda$). In this sense the TDOS is the fundamental spectral object of the theory, while the time-domain memory kernel is its direct transform. The infrared dynamical behavior of the collective coordinate is therefore completely determined by the low-$\lambda$ structure of $\rho(\lambda)$.
@@
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho_{0}\,e^{-\lambda t}.$ (S89)
@@
$K(t)$ $=\rho_{0}\int_{0}^{\Lambda t}\frac{du}{t}\,e^{-u}=\frac{\rho_{0}}{t}\Big(1-e^{-\Lambda t}\Big).$ (S90)
@@
We use the convention $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma^{R}(\omega)$; other sign conventions for the self-energy appear in the literature.
@@
(VII) Conceptual structure (from hidden modes to infrared memory): The derivation above establishes a general structural result. A Markovian dynamical system defined in an extended state space $\{\phi,X_{\lambda}\}$ becomes non-Markovian when the internal degrees of freedom $X_{\lambda}$ are eliminated. The resulting equation for $\phi$ necessarily acquires a causal memory kernel given exactly by the Laplace transform of the time-scale density of states,
@@
$K(t)\simeq\frac{\rho_{0}}{t}.$ (S102)
@@
Appendix D Appendix D: Renormalization-group flow of the Cooper channel in conventional and memory-dominated regimes
@@
Appendix E Appendix E: Finite-temperature regulation of the infrared divergence
@@
Appendix F Appendix F: Spectral function generated by memory–dominated self–energy